# The central engine of low-luminosity radio galaxy 3C 270

### v1

@@ 3.2 Precluding Seyfert I type line profile in the polarized flux spectrum
If the polarized flux contains a Seyfert I-type line profile, can this possibility be robustly ruled out? We adopt the method outlined in Marin et al. (2024) to evaluate what fraction of the genuine type-I profiles would be conspicuous if present in our polarized flux spectrum.
Accurate determination of the polarized continuum flux is essential since the polarization of the continuum is heavily diluted. Using HST imaging polarimetry, where significant polarization has been detected (Capetti et al., 2007), we calculate the polarized flux as a function of aperture radius. The polarized flux flattens at large apertures, yielding an average polarized flux of erg s-1cm-2Å-1. Given that polarized emission lines contribute approximately to this flux, the polarized continuum flux is erg s-1cm-2Å-1, in excellent agreement with our observations.
We scale the real type-I profiles onto the polarized continuum based on their equivalent widths and resample the coadded spectra according to the uncertainties in our polarized flux. The type-I AGN sample is drawn from Schmidt & Green (1983) (23 sources), Stirpe (1990) (5 sources), and Jiang et al. (in prep.; 9 sources). The properties of these sources are summarized in Table A1. This sample comprises highly luminous objects ( mag or log(/erg s-1cm), whose spectra can be compared directly to our polarized flux spectrum due to the absence of host stellar contamination (Shen et al., 2011). We assume that this sample is representative of classified type-I AGNs. We visually inspect the data and find that fewer than 6% (2 out of 37, see Fig.A1) of the type-I profiles would remain undetected.
@@
4 Discussion
Chiaberge et al. (1999) study a sample of FR I radio galaxies using the HST and identify an unresolved optical Compact Central Core (CCC) with a high detection rate, including 3C 270. The presence of a CCC suggests that there is only mild extinction along the line of sight, as substantial extinction would completely obscure the source. Typically, low extinction corresponds to a low column density and further examination on the X-ray emission shows that for most of the sources in the sample, is indeed very low ( cm-2, Balmaverde et al., 2006). However, X-ray studies of 3C 270 require modeling the spectrum with a significantly higher column density of cm-2 (Zezas et al., 2005; Worrall et al., 2010).
The observed discrepancy between high column density and low extinction in 3C 270 could be explained if part of the X-ray obscuring material is not along the line of sight for the optical emission. However, in our favored scenario, we are observing the optical emission through a mirror and do not expect X-ray column density to apply.
5 Conclusion
If the high X-ray column density does mean the optical core is hidden in this object, our data indicate that we can see the hidden nuclear region with a polar mirror. Since we don’t observe broad lines in polarized flux but we do observe continuum and narrow lines, it seems likely that there is no hidden BBB and BLR. Thus, our periscopic view reveals a continuum which is likely beamed synchrotron, and the innermost NLR. The presence of a mid-infrared point source (Asmus et al., 2011, 2014) and warm dust emission (Leipski et al., 2009) in 3C 270 likely arises from the jet or Advection-Dominated Accretion Flow (ADAF) heating the compact torus instead of a BBB. This is our favored hypothesis, making 3C 270 similar to radio-loud AGN Centaurus A (Marin et al. 2024) and radio-quiet AGN NGC 4258 (Barth et al., 1999b). The caveat is that the scattered continuum is only weakly detected, so a commensurate BBB and BLR can’t be totally ruled out.

### v3

@@ 3.2 Precluding Seyfert I type line profile in the polarized flux spectrum
Is it possible to robustly rule out the possibility that the polarized flux contains a Seyfert I-type line profile? We adopt the method outlined in Marin et al. (2025) to evaluate what fraction of the genuine type-I profiles would be conspicuous if present in our polarized flux spectrum.
Accurate determination of the polarized continuum flux is essential since the polarization of the continuum is heavily diluted by starlight in total flux. Using HST imaging polarimetry, where significant polarization has been detected (Capetti et al., 2007), we obtain a polarized flux of erg s-1 cm-2 Å-1 within 1.2′′ radius aperture, comparable to the average seeing of our observations. Given that polarized narrow emission lines contribute approximately 20% to this flux with the same polarization angle, the polarized continuum flux is erg s-1 cm-2 Å-1, in excellent agreement with our observations.
We scale the real type-I profiles onto the polarized continuum based on their equivalent widths and resample the coadded spectra according to the uncertainties in our polarized flux. The type-I AGN sample is drawn from Schmidt & Green (1983) (23 sources), Stirpe (1990) (5 sources), and Jiang et al. (in prep.; 9 sources). The properties of these sources are summarized in Table A1. This sample comprises highly luminous objects ( mag or log(/erg s-1 cm), whose spectra can be compared directly to our polarized flux spectrum due to the absence of host stellar contamination (Shen et al., 2011). We assume that this sample is representative of classified type-I AGNs. We calculate the SNR of the broad wing region in the coadded spectra, defined at the wavelength corresponding to the half-maximum. A broad feature is considered detected if the SNR at this wavelength exceeds 3. We find that fewer than 6% (2 out of 37, see Fig. A1) of the type-I profiles would remain undetected.
@@
4 Discussion and Summary
Chiaberge et al. (1999) studied a sample of FR I radio galaxies using the HST and identified an unresolved optical Compact Central Core (CCC) with a high detection rate, including 3C 270. The presence of a CCC in 3C 270 suggests that there is only mild extinction ( 3) along the line of sight, as substantial extinction would completely obscure the source (Chiaberge et al., 2003). Typically, low extinction corresponds to a low column density and further examination on the X-ray emission shows that for most of the sources in the sample, is indeed very low ( cm-2, Balmaverde et al., 2006). However, X-ray studies of 3C 270 require modeling the spectrum with a significantly higher column density of cm-2 (Zezas et al., 2005; Worrall et al., 2010).
The observed discrepancy between high column density and low extinction in 3C 270 could be explained if most of the X-ray obscuring material is not along the line of sight to the optical emission. In our favored scenario, we are observing the optical emission through a mirror and do not expect X-ray column density to apply.
If the high X-ray column density does mean the optical core is hidden in this object, our data indicate that we can see the hidden nuclear region with a polar mirror. Since we don’t observe broad lines in polarized flux but we do observe continuum and narrow lines, it seems likely that there is no hidden BBB and BLR. Thus, our periscopic view reveals the underlying continuum, including any contributions from Radiatively Inefficient Accretion Flow (RIAF) and beamed synchrotron, and the innermost NLR. With the available data, we can not distinguish which of the two components dominates the scattered continuum. In the simplest FR I – BL Lac unification scheme (Chiaberge et al., 1999; Capetti et al., 2000; Trussoni et al., 2003), a scattered continuum dominated by beamed synchrotron emission would be favored. de Menezes et al. (2020) attempted a broadband SED analysis of 3C 270 and favored a jet-dominated synchrotron self - Compton model over the RIAF model. However, we caution that the optical emission being scattered should be fully considered when performing such analyses, and that multi-wavelength (polarimetric) SED analysis are needed to better distinguish the two components (e.g., Marin, 2018). In either cases, the presence of a mid-infrared point source (Asmus et al., 2011, 2014) and warm dust emission (Leipski et al., 2009) in 3C 270 arising from the BBB heating the compact torus can be ruled out. This is our favored hypothesis, making 3C 270 similar to radio-loud AGN Centaurus A (Marin et al., 2025) and radio-quiet AGN NGC 4258 (Barth et al., 1999c). The caveat is that the scattered continuum is only weakly detected, so a commensurate BBB and BLR can’t be totally ruled out.